Sol Summary – December 4th

MDRS Daily Summary Report for Sol 3

Summary Title: First full up day of activities!

Mission Status: Today really felt like our first day on mars. It was our first full activity day as well as our first day on EVA! Surprisingly it was consensus among our crew that our first full day was actually the easiest day we’ve had so far! I think the excitement of going outside and our team getting into the swing of things has really made us more efficient.

The primary activity for the day was an EVA on all crew to familiarize ourselves with EVA procedures, suit functionality, and operation of the rovers. Many lessons were learned, and it was good that we got the kinks out on a EVA solely suited to it. This will allow our scientific EVAs to be successful later on, as well as ensure that we can safely perform EVAs at a farther distance from the habitat.

Our team is following a specific protocol for depress / repress operations modeled after ISS EVAs. Our team is doing a 25 minute prebreathe in the EVA Prep room to prevent decompression sickness, once complete we are doing a 5 minute depressurization in the actual airlock. During repress we are allocating 5 minutes for repressurization options, though we have the capability to do an ’emergency repress’ if needed at the risk of rupturing crew eardrums.

Below are the several lessons learned we had during EVA.

  1. Peripheral vision is nearly non-existent when wearing the EVA helmets. This makes it difficult to travel together in an ATV convoy as it is impossible to turn around and check on the status of those behind you. We attempted to mitigate this by placing the radios in VOX and keeping in constant voice contact to maintain proper spacing, we also played ‘leapfrog’ with the ATV’s where the leader would stop after several minutes of driving to allow the two behind him to catch up and pass, ensuring no-one was left behind.
  2. We had several communication failures during EVA due to earpieces falling out of ears. We addressed this via using medical tape to secure the earpieces in place.
  3. We had to abort a prebreathe operation due to forgetting to bring supplies into the EVA Prep room. We instituted a supply check prior to closing the Prep Room door. If necessary to break EVA prebreathe protocol we do a 3 to 1 buy back, i.e. if you break prebreathe for 5 minutes, you add 15 minutes to prebreathe to compensate.
  4. Several of our crew members are trained scuba divers. We found that it was difficult to hear other EVA crew while driving the ATV due to engine and wind noise, we instituted a series of hand signals to allow non-verbal communication which quickly and concisely allowed EVA crew status to be conveyed.
  5. Do not wear high pony tails – recommend low braid. A high braid interferes with the way that your head can move within the EVA helmet and vision is confined to roughly a 45 degree angle down and lower.

During the high winds today we heard a bang as the outer airlock door was blown open. The inner airlock door held closed and thus we didn’t all die! We immediately closed the door to the EVA Prep room to form a second seal to the martian atmosphere. During tomorrows familiarization EVA we will enter the EVA Prep room, close the EVA Prep room door, and use that as a depressurization chamber. Immediately upon departure on tomorrows EVA we will repair the outer airlock door.

We also completed our first scientific objectives today, covered below.

Reduced Gravity CPR Experiment: When on Mars there is a possible complication when performing CPR. With your body weight reduced the typical CPR techniques are non functional due to reduced force on the recipient. Our team is practicing techniques under the supervision of our medical officer to allow force application in excess of body weight and measuring output using a springed scaled.

Communication Delay ‘Bricks’ Experiment: Our first two experiment runs of the brick experiment were performed. We started with a 0 time delay baseline, and increased to a 9 minute time delay with verbal information transfer (martian time delay). Our runs will continue to collect more data and determine the most efficient methods of information transfer.

We also completed some major maintenance tasks today

The Musk Observatory has been set up and put through it’s paces and is ready for operation.

Trailer Tank to Static Water Tank transfer operation – Crew completed the transfer this morning completely filling our static water tank. Heartfelt thanks to the Team supporting us for dropping the trailer tank and instructions on how to transfer the water. The MDRS Hab Manual was a roaring success in allowing us to complete this task!

Sol Activity Summary:

  1. Tasks mentioned above
  2. Exercise
  3. Dirt dehydration
  4. Science Dome organization and clean up
  5. Repair of Roof Hatch hole
  6. Our first shower! Heartily enjoying the new heater
  7. Completion of reports (soon)
  1. EVA Request
  2. Sol Summary
  3. Greenhab
  4. Journalist
  5. Operations
  6. Astronomer
  7. Inventory

WiFi range survey Radio check with all radios (all functional except for 1) Look Ahead Plan:

1. Matryoshka EVA #1

2. Communication ‘Lego Bricks’ experiment

3. Sleep Study / crew medical questionnaires

4. Musk Observatory Observation

5. Journalist interviews and filming

6. Medical status surveys with Crew Doctor throughout the day.

7. CPR Experiment

8. HAB Cleaning

9. Radio Range EVA – This activity is cancelled due to 1 EVA per day limit. It will be rescheduled for a future day. We will work with schedulers to ensure they understand this limit for the future.

Anomalies in work:

  1. ATV wobbly wheel
  2. Non Functional Power Strip (removed and replaced)
  3. 3 Non Functional Radios (does not impact operations)
  4. HAB paneling peeling off due to wind damag

Weather: Very mild weather. This was welcome after yesterday.

Crew Physical Status: All crew in good health.

EVA: 2 Familarization EVAs today

Support Requested:

  1. ATV Fuel
  2. Food Drop (not urgent)

.

Astronomy Report – December 4th

Facility Utilized: Musk Observatory

Crewmembers Utilized: Thomas Horn

Supplies Utilized: Solar Telescope / Eyepiece / Camera / Dome

Duration (minutes): 180 minutes

Description of Activity: Radiation is a deadly threat to astronauts both in transit to and on the surface of mars, particularly when the astronauts are exposed to the sun. While astronauts can rely on ground support to identify any solar hazards that develop, communication with the ground cannot be guaranteed. If a Mars Habitat had an independent method of solar observation it could provide critical advance notice that solar flares or other hazards were developing and give direction to crew to shelter in place in radiation storm shelters.

Today’s activity was primarily focused on setting up the solar observatory for operation throughout the mission. A quick survey of the sun was done and no unusual solar activity was spotted. Images below were taken during today’s operation.

Operations Report – December 4th

SOL: SOL 3

Name of person filing report: HuntJ

Non-nominal systems:

  1. Wi-Fi Modem (lack of range).  The wifi in a roughly 5 feet circle from the router, any farther and signal drops off line.  We have tested this and created a signal service map by carrying laptops around the habitat to determine when it picks up signal.  No known fix, will operate internet within this range for now.
  2. HAB Eastern Wall has sustained wind damage, panels are beginning to peel off wall. When wind blows the panels swing and hit support beam causing HAB to vibrate.  In order to fix we need HAB material adhesive delivered, after which we will go EVA and patch the damage,
  3. ATV Yamaha 250 #3 Back Right Wheel is loose, observed to be wobbling during driving.  We will assess a fix during tomorrows EVA – it is possible it simply needs to be tightened, if there is actual mechanical damage then we will move the ATV and declare it off limits.
  4. 3 broken old radios.  All newly delivered radios are functional (thank you supply drop).  Assessing broken radios to determine a fix, if none identified will mark as trash and place in a ‘broken equipment bin’.  4 EVA earpieces were declared nonfunctional, though this may have been due to incorrect settings, will continue to investigate during tomorrows EVA

Generator (hours run): 15.4 Hours

Solar— SOC % : @ 7:34 AM: 65%

                            @ 8:00 AM: 100 % (Generator stopped at this time)

@ 5:34 AM: 42% (Generator started at this time)

Diesel –  48% Full.

Propane –  79 Percent Volume .

Ethanol Free Gasoline (5 Gallon containers for ATV) –   gallons. 2.5 Gallons. One 5 gallon tank empty, second tank half full. Request for extra fuel.

Water (trailer) –  600 Gallons

Water (static) –   500 gallons (We have water again!!)

Trailer to Static Pump used –  Yes

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  35 gallons

Toilet tank emptied: No

ATV’s Used: 3 Yamaha 250’s and 1 Honda ATV used.

Oil Added? No

ATV Fuel Used:  7.5 gallons used to fill all ATVs to full, though the ATVs did not start with full tanks for our mission.

# Hours the ATVs were Used today: 02:30 hours

Notes on ATVs: See above for wobbly ATV wheel issue.

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

 

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Trailer to Static water tank was used today. Pump was dirty and needed cleaning and we did so. Actual process was nominal. We are now full on water again. When inspecting trailer tank water the front tank had a layer of FOD and detritus at the water line. We will keep watch of filter usage in case this increases our burn through rate on filters.

Summary of internet:  Wifi was fixed this morning, though range is limited (see above). Please send hardware manual for the router and modem to allow us to diagnose issues and repair anything that comes up in the future (for example – to allow us to diagnose what certain light combinations signify)

Summary of suits and radios: See above

Summary of Hab operations: Brick Communication Experiment / CPR experiments successful, exercise on all crew. EVA Outer Door repaired.

Summary of GreenHab operations: No fire extinguisher observed – please deliver in future supply drop.

Summary of ScienceDome operations: Soil Salinity experiments (previous crew), dehydrating dirt for use as counterbalance and weights in exercise equipment (filling bags and medicine balls)

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support:

Need extra fuel for ATVs. Of the two red fuel cans, one is empty, one is half full.

Pre-Mission food inventory is complete, will be sent as it’s own report. Please reference that report for requested supplies to address gaps in current inventory. Please give ETA for supply drop so we can adequately ration food.

Thank you for your assistance!

GreenHab Report – December 3rd

GreenHab Report

Trisha Randazzo

December 3, 2017

Sol 2

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

·         Windy so unable to keep door open

Heating

·         Functioning nominally

Cooling

·         Not availabloe

Both heating and cooling

·         Nominal

Shade cloth on/off

·         On

Average temperature: (N/A until new sensor is delivered)

·         N/A

Changes to crops:  Note all emergence of seedlings, death of seedlings, etc

·         One tomato is doing poorly, but everyone else is growing quite well!!

Daily water usage for crops:

·         Complete with no issues

Time(s) of watering for crops:

·         12:55

Morning research observations: (Currently not operational)

·         N/A

Changes to research plants:

·         N/A

Daily watering and amount of water used:

·         2 gallons of all soils

Aquaponics:  (Currently not operational)

·         N/A

Narrative:  Any other information you want to share

·         Planted crew gifts of paper-whites. Two containers in the Green-Hab, one container in the Hab.

Support/supplies needed:

·         Awaiting:

Operations Report – December 3rd

Crew 184 Operations Report  03/12/2017

SOL: SOL 2

Name of person filing report: HuntJ

Non-nominal systems: Toilet Tank, Water (static)

Notes on non-nominal systems:  Toilet Tank: Tank was particularly filled up during purge and caused all crew to have to move upstairs for duration of purge. Workaround is to purge every other day of ops (e.g Sol 4 is next purge) to make upkeep easier on all.

We also request a resupply of the TB-Cide Quat Cleaner Deordorizer/Disinfectant. We are out after the last purge.

 

                                                    Water (Static): Water is at 50 Gallons which is alarmingly low. Our trailer water is at 1000 gallons but we need the means to transfer the water. Request to get Trailer water moved to Static water or permission and high level instructions to do it ourselves.

Generator (hours run): 13.2 Hours

Solar— SOC % (Before generator is run at night) Not nightfall as of submission, but will report this info in the next report.

Diesel –  50% full

Propane –  Did not check (No EVA), assume to be 79 percent volume due to past readings and trends.

Ethanol Free Gasoline (5 Gallon containers for ATV) –   gallons. 5 gallons

Water (trailer) –   1000 Gallons

Water (static) –   50 Gallons

Trailer to Static Pump used –  No

Water (loft) – Static to Loft Pump used –  Yes

Water Meter:  20 gallons (46 once Pump used)

Toilet tank emptied: Yes

ATV’s Used: None

Oil Added? No

ATV Fuel Used:  None

# Hours the ATVs were Used today: 00

Notes on ATVs: N/A

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments:

Summary of internet:  Still good so far.

Summary of suits and radios: All nominal

Summary of Hab operations: Static Water Tank almost empty, need means to transfer Trailer water over to Static.

Summary of GreenHab operations:  All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Is there a way we can check to see how much internet we are using during the day? We tried the link provided in a different log template but it didn’t take us to anything.

Sol Summary – December 3rd

Crew 184 Daily Summary Report  03 Dec 2017

MDRS Daily Summary Report for sol 2

Summary Title:  Getting settled at MDRS!

Mission Status:  Today has been a busy day preparing both ourselves and the MDRS for the rest of our mission.  We spoiled ourselves with a hearty breakfast today (and to get rid of some of our fresh food!).

Today has been one of excitement preparing for our first EVA tomorrow.  We have encountered our first mission failure!

During the high winds today we heard a bang as the outer airlock door was blown open.  The inner airlock door held closed and thus we didn’t all die!  We immediately closed the door to the EVA Prep room to form a second seal to the martian atmosphere.  During tomorrows familiarization EVA we will enter the EVA Prep room, close the EVA Prep room door, and use that as a depressurization chamber.  Immediately upon departure on tomorrows EVA we will repair the outer airlock door.

We have discovered today that all the planning in the world still requires real-time flexibility as numerous activities took much longer than we expected.  For example the Exercise Bike took several hours and modifications to assemble and get working correctly, much more than we had allocated.  That being said, we have successfully completed all objectives and are ready to begin completing our science objectives tomorrow!.

Sol Activity Summary:

1.    Inventoried and organized the MDRS Hab and Greenhab

2.    Completed all reports

a.    Greenhab

b.    HSO

c.    Mission Plan

d.    Operations Report

e.    Sol Summary

f.     EVA Request

g.    Journalist

3.    Assembled exercise equipment (Stationary Bike / Weights)

4.    Got more detailed training from previous crewmembers

5.    Crew Exercise

6.    Internet troubleshooting (WiFi modem nonfunctional – connection working straight to modem)

7.    Cleaning and organizing MDRS

8.    Radio check with all radios (all functional except for 1)

9.    Communication check at all MDRS buildings with Habitat

10. Neuroblock Feasability Training

Look Ahead Plan:

1.    Familiarization EVA

2.    Communication ‘Lego Bricks’ experiment

3.    Sleep Study / crew medical questionnaires

4.    Musk Observatory Training and Observation

5.    Journalist interviews and filming

6.    Medical status surveys with Crew Doctor throughout the day.

7.    CPR Training

a.    Practice CPR techniques with crew medical doctor on CPR medical dummy

See schedule for tomorrow attached to this email.

Anomalies in work:

8.    Non Functional WiFi Modem

9.    Non Functional Power Strip (removed and replaced)

10. 1 Non Functional Radio

11. EVA Airlock Outer Hatch Failure (failed open) due to high winds.

12. Roof Hatch blown open continually (will close when winds die down)

Weather:  Very windy!  We have had two wind related MDRS failures.

Crew Physical Status:  All crew in good health.

EVA: No EVA today

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Sol Summary

Mission Plan

Support Requested:

1.    We can’t find a template for Daily HSO Report.  Please provide what is desired in this report.

2.    Water Transfer from Delivery Tank to Static Tank

3.    More TB-Cide Quat Cleaner (we are out)

4.    Recommendation for wi-fi router failure.

a.    We have reset the router, checked our allocation, restarted laptop, and router is non-functional.  When hooking up laptop directly to modem internet works well.

Greenhab Report – December 3rd

GreenHab Report

Trisha Randazzo

December 3, 2017

Sol 2

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

· Windy so unable to keep door open

Heating

· Functioning nominally

Cooling

· Not availabloe

Both heating and cooling

· Nominal

Shade cloth on/off

· On

Average temperature: (N/A until new sensor is delivered)

· N/A

Changes to crops: Note all emergence of seedlings, death of seedlings, etc

· One tomato is doing poorly, but everyone else is growing quite well!!

Daily water usage for crops:

· Complete with no issues

Time(s) of watering for crops:

· 12:55

Morning research observations: (Currently not operational)

· N/A

Changes to research plants:

· N/A

Daily watering and amount of water used:

· 2 gallons of all soils

Aquaponics: (Currently not operational)

· N/A

Narrative: Any other information you want to share

· Planted crew gifts of paper-whites. Two containers in the Green-Hab, one container in the Hab.

Support/supplies needed:

· Awaiting:

Sol Summary – December 2nd

Sol 1

Summary: MDRS Training Day!

Author: Tom Horn

Mission Status:  Mission is starting great and we are enjoying our first day on Mars! Today was a busy day as we spent the day out of sim getting training from the previous crew and Shannon. We are excited for our first day ‘in sim’ tomorrow. We are taking a ‘light duty’ day tomorrow as we focus on finishing move in to the hab and getting acclimated. Due to this we are NOT doing an EVA tomorrow.

Notes for CAPCOM. Please communicate with our crew using this email address, mdrscrew184@gmail.com. This ensures our entire crew sees all communications. Please add this email address to the mission support distribution so in the future we can email these reports to you from that address.

One of the major research goals for our team is to follow a ‘Martian Day’ of 24:40 and study the effects on our team from following a longer day. Our team is going to follow the schedule below. We understand the the ‘CAPCOM Communication Window’ is between 7-9pm, we will always be awake during this time and will ensure to get reporting completed during this window, including any and all EVA requests. Note that we won’t be following a normal day so if a request is sent to our crew during our ‘sleep time’ our response will be delayed.

Further details will follow in the Mission Plan to be sent tomorrow.

Flight Day Calendar Day Wake Bed Twilight Start Sunrise Sunset Twilight End
Day 1 12/2 Not Scheduled 8:00 PM 6:53:57 AM 7:23:08 AM 5:02:24 PM 5:31:34 PM
Day 2 12/3 4:00 AM 8:40 PM 6:54:50 AM 7:24:03 AM 5:02:17 PM 5:31:30 PM
Day 3 12/4 4:40 AM 9:20 PM 6:55:41 AM 7:24:57 AM 5:02:13 PM 5:31:28 PM
Day 4 12/5 5:20 AM 10:00 PM 6:56:32 AM 7:25:50 AM 5:02:11 PM 5:31:28 PM
Day 5 12/6 6:00 AM 10:40 PM 6:57:21 AM 7:26:41 AM 5:02:10 PM 5:31:30 PM
Day 6 12/7 6:40 AM 11:20 PM 6:58:10 AM 7:27:32 AM 5:02:12 PM 5:31:34 PM
Day 7 12/8 7:20 AM 12:00 AM 6:58:57 AM 7:28:22 AM 5:02:16 PM 5:31:41 PM
Day 8 12/9 8:00 AM 12:40 AM 6:59:44 AM 7:29:10 AM 5:02:23 PM 5:31:49 PM
Day 9 12/10 8:40 AM 1:20 AM 7:00:29 AM 7:29:56 AM 5:02:31 PM 5:31:59 PM
Day 10 12/11 9:20 AM 2:00 AM 7:01:13 AM 7:30:42 AM 5:02:42 PM 5:32:11 PM
Day 11 12/12 10:00 AM 2:40 AM 7:01:55 AM 7:31:26 AM 5:02:54 PM 5:32:25 PM
Day 12 12/13 10:40 AM 3:20 AM 7:02:36 AM 7:32:08 AM 5:03:09 PM 5:32:41 PM
Day 13 12/14 11:20 AM 4:00 AM 7:03:16 AM 7:32:49 AM 5:03:26 PM 5:32:58 PM
Day 14 12/15 12:00 PM 4:40 AM 7:03:55 AM 7:33:29 AM 5:03:44 PM 5:33:18 PM
Day 15 12/16 12:40 PM Not Scheduled 7:04:32 AM 7:34:06 AM 5:04:05 PM 5:33:40 PM
Day 16 12/17 Not Scheduled – Handover Day 7:05:07 AM 7:34:42 AM 5:04:28 PM 5:34:03 PM

 

Sol Activity Summary: We arrived at 12pm today and spent the rest of the day learning about the MDRS with Shannon. We enjoyed our last ‘out of sim’ dinner at Stan’s burger shack with Shannon and the previous crew.

Look Ahead Plan:

Tomorrow our primary tasks are:

  1. Finishing move in to the HAB
  2. Completing an MDRS Inventory of supplies (food and other)
  3. Finish cleaning MDRS to start the mission.
  4. ScienceDome / Greenhab / Solar Observatory Detailed Inspection
  5. Completion of required reporting
    1. Mission Plan
    2. Health and Safety Officer Report
    3. If other ‘beginning of sim’ reports are requested, please let us know!

Anomalies in work:  None

Weather: Unseasonably warm!

Crew Physical Status: All crew in good physical health. Brandon Ferguson from the previous crew is staying with our crew until Monday night.

EVA: No EVA

Reports to be filed:

  1. Sol Summary (this is intended to include the commanders report as a narrative included in Mission Status. If this did not hit sufficient detail for this please let me know what details are desired and I will include them in future reports)
  2. Operational Report

Support Requested: None

Journalist Report & Crew Photos – December 2nd

Crew 184 Journalist Report

Willie Schumann

02 December 2017

Title Day One on Mars

Narrative We are on Mars. It’s unbelievable, but yet so real. After half year of planning , we arrived yesterday to the habitat and it was a rush of first impressions. We cleaned some larger items of the previous crews and made ourselves at home. The effects of the long Martian day, which is 40 minutes longer then a day on earth, kicked in and we were quite tired. We dropped into our beds happy and exhausted, which is the best state of mind when you finish a day.

There is always beauty in new beginnings, and so was the morning of our first Martian Day. We got up really early. There is still so much to be done before we can start our research and breath life into the plentiful projects of crew 184. Nonetheless, we spoiled ourselves with a hearty breakfast of scrambled eggs. We are trying to use our fresh food as soon as possible. Especially in the first days we want to keep moral high, as the six of us have to get used to the new environment.

After breakfast, our second in command Trisha Randazzo gave everyone in the crew a little bag of presents. What a surprise, which could only be topped by Mars itself. The first sunrise on the red planet blew our minds. The colors of the distant sun, leaping over the Martian hills was something I had never seen on Earth. Maybe the euphoria tainted my perception, but anyway, it was darn beautiful.

Even though we were not assigned on an EVA, just yet, the day was eventful. First Officer Randazzo and Crew Engineer Hunt started assembling the workout bicycle really early and had a steep learning curve. It’s seems to be a universal truth, that assembling is as hard on Mars as it is on Earth, and that improvisation is always king. It’s not important how you get there, but that is works and that you gained knowledge in the process.

It’s not a surprise that the crew meals on this first day have been a very bonding experience. I hope we can keep this energy, even if we run out of fresh food and of inspiration how to prepare it. We cherish being here together and looking forward for the days to come. We planted six paper white flower in the green hab this noon, one for every crewmember. So we set already the first seeds of new life on this strange planet. We try to keep them alive. If we succeed, we will probably manage most other challenges that wait for us.

Tomorrow we will head out for our first EVA’s. I think, that is when the magic really starts. We are familiar with our new home now, but now we want to see the land, that surrounds us. We hope we can manage to wear the new space suits without larger problems, as we have heard, that you have to be fit. That’s another reason, why we plan an intense workout regime during our stay. I think, even if we struggle in the beginning of our mission, practice will show its effect and we can start to enjoy our exploration of Mars to the fullest.

Personal Logbook As a Journalist and filmmaker I have certain goals like everyone else of my crewmates. I am very excited about the time to come and conscious about my tasks and the things I need to do, to create great stories. While today was mainly spent with preparations and adjustments, tomorrow will be opening a great window of opportunity for fantastic pictures, when we will finally go outside.

I have filmed a lot today, but I think, that the scenes, that follow, will be more meaningful. As now, I haven’t managed to make to many interviews and it will be a challenge to find new corners and angles in the habitat to film. I was a little tired this morning, but worked on autopilot through most of the day. Patience and motivation will be important to make the most of my time on Mars. I guess, these are qualities that are essential for every astronaut and storyteller.

Willie Schumann, Journalist, Crew 184

Operations Report – December 2nd

Crew 184 Operations Report 02/12/2017

SOL: SOL 1

Name of person filing report: HuntJ

Non-nominal systems: All Nominal

Notes on non-nominal systems: N/A

Generator (hours run): N/A (Arrived during day when already off)

Solar— SOC % (Before generator is run at night) 80%

Diesel – 50% Full.

Propane – 80 Percent Volume .

Ethanol Free Gasoline (5 Gallon containers for ATV) – gallons. 5 gallons

Water (trailer) – 1000 Gallons!!!.

Water (static) – 500 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – No

Water Meter: 46 gallons

Toilet tank emptied: No

ATV’s Used: Honda, 300, 350.1, 350.2, 350.3

Oil Added? No

ATV Fuel Used: Negligible (Test drive only)

# Hours the ATVs were Used today: 00:20 hours

Notes on ATVs: They’re running great!!

Deimos rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where? No

General notes and comments: Happy to see the Red Planet!

Summary of internet: So far so good!

Summary of suits and radios: All nominal

Summary of Hab operations: All nominal

Summary of GreenHab operations: All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: All nominal

Questions, concerns and requests to Mission Support: Everything is Awesome!!